2. The average starting salary for this year's graduates at a large university (LU) is $50,000 with a standard deviation of $8,000. Furthermore, it is known that the starting salaries are normally distributed.

a. What is the probability that a randomly selected LU graduate will have a starting salary of at least $61,900?

b. Individuals with starting salaries of less than $36,800 receive a low income tax break. What percentage of the graduates will receive the tax break?

Please if it can be typed so I can better view how the answer was solved
Answer #1

\\\text{Let X denotes the starting salary of a random selected LU graduate}\\ \text{Given }X\sim N(50000,8000^2) \\(a)\text{Probability that a randomly selected LU graduate will have atleast 61900 as starting salary} \\=P(X<61900)\\=P(\frac{X-50000}{8000}<\frac{61900-50000}{8000}) \\=P(Z<19/80) \\=\Phi(19/80) \\=0.5938655 \\(b)\text{Proportion of individuals that receive tax break} \\=P(X<36800) \\=P(\frac{X-50000}{8000}<\frac{36800-50000}{8000}) \\=P(Z<-1.65) \\=0.04947147